2. calculate energy intake for a 4 month female infant with a weight of 6 kg taking 7 oz at 8 am,12 pm, 4 pm, 8 pm, and 2 am. The infant is consuming 20 kcal/oz formula.

a) explain how you did it by showing your calculations

b)calculate both in total calories for the day and in kcal/kg/d

c)assess if recommendations are below, adequate or above recommendations
Answer #1

Infant have taken 7 oz 5 times

= =7* 5

=35 oz

1 oz=30 ml

= 35*30

=1050 ml

C: acc to who recommendation for infant age 0-4 month fluid requirement is 750 ml per day.

so it’s above recommendations.

b. 20 kcal per oz

Total kcal in a day =20*35

Total calories in day =700 kcal

total weight=6 kg

Per kg calories taken = 700/6

= 117 kcal/ kg/day